### v1

# Computational complexity of the recoverable robust shortest path problem in acyclic digraphs

Abstract
In this paper, the recoverable robust shortest path problem in acyclic digraphs is considered. The interval budgeted uncertainty representation is used to model the uncertain second-stage costs. The computational complexity of this problem has been open to date. In this paper, we prove that the problem is strongly NP-hard even for the case of layered acyclic digraphs. We also show that for the discrete budgeted uncertainty, the problem is not approximable unless P=NP.
Keywords: robust optimization, interval data, recovery, shortest path, computational complexity
1 Problem formulation
We are given a directed acyclic graph $G=(V,A)$, where $|V|=n$ and $|A|=m$. Two nodes in $V$ are distinguished as a source $s\in V$ and a sink $t\in V$. Let $\Phi$ be the set of all $s$-$t$ paths in $G$ ($X\in\Phi$ is the set of arcs forming the path). For each arc $e\in A$, a first-stage cost $C_{e}\geq 0$ is specified. The second-stage arc costs are uncertain. We assume they belong to an uncertainty set $\mathcal{U}\subseteq\mathbb{R}_{+}^{m}$. In the recoverable robust shortest path problem, we compute a first-stage path $X\in\Phi$. Then, after a second-stage cost scenario $\boldsymbol{c}\in\mathcal{U}$ is revealed, the path $X$ can be modified to some extent. That is, we can choose a second-stage path $Y$ in some prescribed neighborhood $\Phi(X,k)$ of $X$, where $k\geq 0$ is a given recovery parameter. We consider three types of neighborhoods, called arc-inclusion, arc-exclusion, and arc-symmetric difference, respectively:
$\Phi^{\mathrm{incl}}(X,k)$ =$=\{Y\in\Phi\,:\,|Y\setminus X|\leq k\},$ (1)
$\Phi^{\mathrm{excl}}(X,k)$ =$=\{Y\in\Phi\,:\,|X\setminus Y|\leq k\},$ (2)
$\Phi^{\mathrm{sym}}(X,k)$ =$=\{Y\in\Phi\,:\,|(Y\setminus X)\cup(X\setminus Y)|\leq k\}.$ (3)
The recoverable robust shortest path problem is defined as follows:
Rec Rob $\textsc{Rec Rob SP}:\;\min_{X\in\Phi}\left(\sum_{e\in X}C_{e}+\max_{\boldsymbol{c}\in\mathcal{U}}\min_{Y\in\Phi(X,k)}\sum_{e\in Y}c_{e}\right).$ (4)
In this paper, we use the interval uncertainty representation for the uncertain second-stage arc costs. That is, for each arc $e\in A$ we specify a nominal second-stage cost $\hat{c}_{e}$ and the maximum deviation $\Delta_{e}$ of the second-stage cost from its nominal value. Hence, $c_{e}\in[\hat{c}_{e},\hat{c}_{e}+\Delta_{e}]$, $e\in A$, and
$\mathcal{U}=\{\boldsymbol{c}\in\mathbb{R}_{+}^{m}\,:\,c_{e}\in[\hat{c}_{e},\hat{c}_{e}+\Delta_{e}],e\in A\}.$ (5)
To control the amount of uncertainty, we consider two budgeted versions of $\mathcal{U}$, namely the discrete [2] and continuous [10] budgeted uncertainty sets, defined as follows:
$\mathcal{U}(\Gamma^{d})=\{\boldsymbol{c}\in\mathcal{U}\,:\,|\{e\in A\,:\,c_{e}>\hat{c}_{e}\}|\leq\Gamma^{d}\},$ (6)
$\mathcal{U}(\Gamma^{c})=\{\boldsymbol{c}\in\mathcal{U}\,:\,\sum_{e\in A}(c_{e}-\hat{c}_{e})\leq\Gamma^{c}\}.$ (7)
The parameters $\Gamma^{d}$ and $\Gamma^{c}$ are called budgets. The discrete budget $\Gamma^{d}\in\{0,\dots,m\}$ denotes the maximum number of the second-stage costs that can differ from their nominal values. On the other hand, the continuous budget $\Gamma^{c}\geq 0$ denotes the maximum total deviation of the second-stage costs from their nominal values. Both uncertainty sets are commonly used in robust optimization as they allow us to reduce the prices of robustness of the solutions computed. They reduce to $\mathcal{U}$ when $\Gamma^{c}=\Gamma^{d}=0$.
The concept of recoverable robustness was first proposed in [9]. The recoverable robust version of the shortest path problem was first investigated in [3, 4]. The problem is known to be strongly NP-hard for general digraphs and the interval uncertainty set $\mathcal{U}$. This hardness result remains true for all three types of neighborhoods (1)-(3) (see,  [4, 10, 5, 7]). When the input graph $G$ is acyclic, the problem can be solved in polynomial time for the uncertainty set $\mathcal{U}$ and all neighborhoods (1)-(3). This result has recently been shown in [8]. However, the complexity status of the problem for acyclic digraphs and the budgeted uncertainty sets $\mathcal{U}(\Gamma^{d})$ and $\mathcal{U}(\Gamma^{c})$ has been open to date. In the next section, we will show that adding budgets to $\mathcal{U}$ makes the problem strongly NP-hard.
2 Computational complexity results
The idea of the reductions proposed in this section is similar to that in [1] for the min-max regret shortest path problem. Consider the following Hamiltonian Path problem, which is known to be strongly NP-complete (see, e.g., [6]).
Hamitonian Path
Input: Given a directed graph $G=(V,A)$.
Question: Does $G$ have a Hamiltonian path, that is, a simple directed path that visits all nodes of $G$ exactly once?
Theorem 1.
The Rec Rob SP problem in acyclic digraphs, uncertainty set $\mathcal{U}(\Gamma^{c})$, any neighborhood in (1)-(3) is strongly NP-hard, even for $\Gamma^{c}=1$.
Proof.
Let $G=(V,A)$, $|V|=n$, $|A|=m$, be an instance of Hamiltonian Path. We build the corresponding instance of Rec Rob SP as follows. The graph $G^{\prime}$ has $2n$ copies of the nodes in $V=\{v_{1},\dots,v_{n}\}$, labeled as $v_{i}^{j}$, $i\in[n]$, $j\in[2n]$, and two additional nodes denoted by $s$ and $t$. The arcs of $G^{\prime}$ are formed as follows. The vertical arcs connect $v_{i}^{j}$ with $v_{i}^{j+1}$ for $i\in[n]$, $j\in[2n-1]$, node $s$ with $v_{i}^{1}$, $i\in[n]$, and nodes $v_{i}^{2n}$ with $t$, $i\in[n]$. The vertical arcs form $n$ disjoint vertical $s$-$t$ paths $P_{1},\dots,P_{n}$ in $G^{\prime}$, all of which have exactly $12n+$ 1 arcs. The diagonal arcs connect $v_{i}^{2j}$ with $v_{k}^{2j+1}$ for $(v_{i},v_{k})\in A$, $j\in[n-1]$ (the graph $G^{\prime}$ is constructed in the same way as in [1, the proof of Theorem 2], where the min-max regret shortest path problem under the interval uncertainty is considered). A sample construction is shown in Figure 1. Set any positive budget $\Gamma^{c}>0$, say $\Gamma^{c}=1$, and the recovery parameter $k=2n$ for arc-inclusion (arc-exclusion neighborhood) and $k=4n$ for the arc-symmetric difference neighborhood. All arcs in $G^{\prime}$ have first-stage costs equal to 0. The second-stage costs of the vertical arcs are in the interval $[0,M]$, and the second-stage costs of the diagonal arcs are in the interval $[M,M]$, where $M$ is any constant greater than $\Gamma^{c}$. We will show that $G$ has a Hamiltonian path if and only if there is a solution to Rec Rob SP in $G^{\prime}$ with the total cost at most $\frac{\Gamma^{c}}{n}$.
Figure 1: A sample digraph $G$ and the corresponding digraph $G^{\prime}$.
$\Rightarrow$ Let $P$ be a Hamiltonian path in $G$ that visits the nodes of $G$ in order $v_{i_{1}},\dots,v_{i_{n}}$. Consider path $X$ in $G^{\prime}$ that contains the arcs
$(s,v_{i_{1}}^{1}),(v_{i_{1}}^{1},v_{i_{1}}^{2}),(v_{i_{1}}^{2},v_{i_{2}}^{3}),(v_{i_{2}}^{3},v_{i_{2}}^{4}),(v_{i_{2}}^{4},v_{i_{3}}^{5}),\dots(v_{i_{n}}^{2n},t)$
These arcs exist by the construction of $G^{\prime}$. An example is shown in Figure 1. The Hamiltonian path in $G$, which visits the nodes of $G$ in order $v_{1},v_{3},v_{2},v_{4}$, corresponds to the path composed of the arcs shown in red on the corresponding graph $G^{\prime}$. Observe that $X$ uses at least one arc from each vertical path $P_{1},\dots,P_{n}$. Therefore, each vertical path $P_{i}$ is in $\Phi(X,k)$ and can be used as the second-stage (recovery) path. In the worst second-stage cost scenario for $X$, the budget $\Gamma^{c}$ is shared equally among the arcs of all vertical paths (all vertical paths must have the same second-stage cost). Because the first-stage cost of $X$ is 0 and the second-stage cost of the best second-stage path is $\frac{\Gamma^{c}}{n}$, the total cost of the solution to Rec Rob SP is $\frac{\Gamma^{c}}{n}$.
$\Leftarrow$ Assume that there is a solution to Rec Rob SP with a total cost of at most $\frac{\Gamma^{c}}{n}$. Let $X$ be the first-stage path in this solution. We claim that $X$ must use at least one arc from each vertical path $P_{1},\dots,P_{n}$. Indeed, suppose that one of these paths, say $P_{l}$, does not intersect with $X$. Because $k=2n$, the path $P_{l}\notin\Phi(X,k)$. Since the second-stage costs of the diagonal arcs are equal to $M\geq\Gamma^{c}$, no diagonal arc can be used in the second-stage path. Therefore, in a worst scenario for $X$, the budget $\Gamma^{c}$ is not allocated to the arcs in path $P_{l}$. Consider scenario $\boldsymbol{c}$ in which $\Gamma^{c}$ is equally shared to all the arcs of the remaining $n-1$ vertical paths. In what follows, every second-stage path in $\Phi(X,k)$ has a cost of at least $\frac{\Gamma^{c}}{n-1}$, which contradicts the assumption that there is a solution to Rec Rob SP with the total cost of at most $\frac{\Gamma^{c}}{n}$. We can now retrieve a Hamiltonian path in $G$ from the diagonal arcs in $X$. Notice first that $X$ has exactly $12n+$ 1 arcs ($G^{\prime}$ is a layered digraph) and contains exactly $n-1$ diagonal arcs. Indeed, the diagonal arcs must connect $n$ disjoint vertical paths, so the number of diagonal arcs is at least $n-1$. On the other hand, the number of diagonal arcs cannot be greater than $n-1$, since the number of arcs in $X$ is equal to $12n+$ 1 (we need $n$ arcs from the vertical paths and 2 arcs connecting $s$ and $t$ in $X$). Now, it is easy to see that the diagonal arcs in $X$ correspond to the arcs on a Hamiltonian path in $G$. ∎
Theorem 2.
[…]
Proof.
Given an instance $G$ of Hamiltonian Path, the construction of $G^{\prime}$ is the same as in the proof of Theorem 1. We again set the recovery parameter $k=2n$ for the arc-inclusion (arc-exclusion) neighborhood and $k=4n$ for the arc-symmetric difference neighborhood. The first-stage costs of all arcs in $G^{\prime}$ are 0. The second-stage costs of all vertical arcs in $G^{\prime}$ are in the interval $[0,0]$, except for the dashed arcs $(s,v_{i}^{1})$, $i\in[n]$, whose second-stage costs are in the interval $[0,1]$ (see the dashed arcs in Figure 1). The second-stage costs of all diagonal arcs are in the interval $[1,1]$. Let $\Gamma^{d}=n-1$. We claim that there is a Hamiltonian path in $G$ if and only if there is a solution to Rec Rob SP in $G^{\prime}$ with the cost equal to 0. The proof of this claim is based on the same reasoning as in the proof of Theorem 1. Graph $G$ has a Hamiltonian path if and only if there is a first-stage path $X$ in $G^{\prime}$ that uses at least one arc from each vertical path $P_{1},\dots,P_{n}$. Because, in the worst second-stage cost scenario, we can set the cost of at most $n-1$ dashed arcs to 1, $G$ has a Hamiltonian path if and only if there is a second-stage path in $\Phi(X,k)$ with the cost equal to 0. Using the fact that the first-stage cost of each solution equals 0, we are done. ∎
Acknowledgements
The authors were supported by the National Science Centre, Poland, grant 2022/45/B/HS4/00355.