### v2

# Computational complexity of the recoverable robust shortest path problem in acyclic digraphs

Abstract
In this paper, the recoverable robust shortest path problem in acyclic digraphs is considered. The interval budgeted uncertainty representation is used to model the uncertain second-stage costs. In this paper, we prove that the problem is strongly NP-hard even for the case of layered acyclic digraphs and it is not approximable within $2^{\log^{1-\epsilon}n}$ for any $\epsilon>0$, unless NP $\subseteq$ DTIME$(n^{\mathrm{poly}\log n})$ in general acyclic digraphs for the continuous budgeted uncertainty. We also show that for the discrete budgeted uncertainty, the problem is not approximable unless P=NP in layered acyclic digraphs.
Keywords: robust optimization, interval data, recovery, shortest path, computational complexity
1 Problem formulation
We are given a directed acyclic graph $G=(V,A)$, where $|V|=n$ and $|A|=m$. Two nodes in $V$ are distinguished as a source $s\in V$ and a sink $t\in V$. Let $\Phi$ be the set of all $s$-$t$ paths in $G$ ($X\in\Phi$ is the set of arcs forming the path). For each arc $e\in A$, a first-stage cost $C_{e}\geq 0$ is specified. The second-stage arc costs are uncertain. We assume they belong to an uncertainty set $\mathcal{U}\subseteq\mathbb{R}_{+}^{m}$. In the recoverable robust shortest path problem, we compute a first-stage path $X\in\Phi$. Then, after a second-stage cost scenario $\boldsymbol{c}\in\mathcal{U}$ is revealed, the path $X$ can be modified to some extent. That is, we can choose a second-stage path $Y$ in some prescribed neighborhood $\Phi(X,k)$ of $X$, where $k\geq 0$ is a given recovery parameter. We consider three types of neighborhoods, called arc-inclusion, arc-exclusion, and arc-symmetric difference, respectively:
$\Phi^{\mathrm{incl}}(X,k)$ =$=\{Y\in\Phi\,:\,|Y\setminus X|\leq k\},$ (1)
$\Phi^{\mathrm{excl}}(X,k)$ =$=\{Y\in\Phi\,:\,|X\setminus Y|\leq k\},$ (2)
$\Phi^{\mathrm{sym}}(X,k)$ =$=\{Y\in\Phi\,:\,|(Y\setminus X)\cup(X\setminus Y)|\leq k\}.$ (3)
The recoverable robust shortest path problem is defined as follows:
Rec Rob $\textsc{Rec Rob SP}:\;OPT=\min_{X\in\Phi}\left(\sum_{e\in X}C_{e}+\max_{\boldsymbol{c}\in\mathcal{U}}\min_{Y\in\Phi(X,k)}\sum_{e\in Y}c_{e}\right).$ (4)
In this paper, we use the interval uncertainty representation for the uncertain second-stage arc costs. That is, for each arc $e\in A$ we specify a nominal second-stage cost $\hat{c}_{e}$ and the maximum deviation $\Delta_{e}$ of the second-stage cost from its nominal value. Hence, $c_{e}\in[\hat{c}_{e},\hat{c}_{e}+\Delta_{e}]$, $e\in A$, and
$\mathcal{U}=\{\boldsymbol{c}\in\mathbb{R}_{+}^{m}\,:\,c_{e}\in[\hat{c}_{e},\hat{c}_{e}+\Delta_{e}],e\in A\}.$ (5)
To control the amount of uncertainty, we consider two budgeted versions of $\mathcal{U}$, namely the discrete [3] and continuous [12] budgeted uncertainty sets, defined as follows:
$\mathcal{U}(\Gamma^{d})=\{\boldsymbol{c}\in\mathcal{U}\,:\,|\{e\in A\,:\,c_{e}>\hat{c}_{e}\}|\leq\Gamma^{d}\},$ (6)
$\mathcal{U}(\Gamma^{c})=\{\boldsymbol{c}\in\mathcal{U}\,:\,\sum_{e\in A}(c_{e}-\hat{c}_{e})\leq\Gamma^{c}\}.$ (7)
The parameters $\Gamma^{d}$ and $\Gamma^{c}$ are called budgets. The discrete budget $\Gamma^{d}\in\{0,\dots,m\}$ denotes the maximum number of the second-stage costs that can differ from their nominal values. On the other hand, the continuous budget $\Gamma^{c}\geq 0$ denotes the maximum total deviation of the second-stage costs from their nominal values. Both uncertainty sets are commonly used in robust optimization as they allow us to reduce the price of robustness of the solutions computed. They reduce to $\mathcal{U}$ when $\Gamma^{c}=\Gamma^{d}=0$.
The concept of recoverable robustness was first proposed in [11]. The recoverable robust version of the shortest path problem was first investigated in [4, 5]. The problem is known to be strongly NP-hard for general digraphs and the interval uncertainty set $\mathcal{U}$. This hardness result remains true for all three types of neighborhoods (1)-(3) (see,  [5, 12, 6, 10]). When the input graph $G$ is acyclic, the problem can be solved in polynomial time for the uncertainty set $\mathcal{U}$ and all neighborhoods (1)-(3). This result has recently been shown in [10]. It turns out that adding budgets to $\mathcal{U}$ increases the computational complexity of the problem. Indeed, it has been shown that the problem is weakly NP-hard even for a very restricted class of digraphs, namely, arc series-parallel ones [8]. In this paper, we show that the problem becomes substantially more challenging for more general digraphs.
2 Computational complexity results
Consider continuous budgeted uncertainty.
Theorem 1.
The Rec Rob SP problem is strongly NP-hard in layered digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{c})$, if $k$ is part of the input.
Proof.
We will show a reduction from the Hamiltonian Path problem, which is known to be strongly NP-complete (see, e.g., [7]).
Hamiltonian Path
Input: Given a directed graph $G=(V,A)$.
Question: Does $G$ have a Hamiltonian path, that is, a simple directed path that visits all nodes of $G$ exactly once?
Let $G=(V,A)$, $|V|=n$, $|A|=m$, be an instance of Hamiltonian Path. We build the corresponding instance of Rec Rob SP as follows. A graph $G^{\prime}$ is constructed in the same way as in [2, the proof of Theorem 2], where the min-max regret shortest path problem under the interval uncertainty is considered. It has $2n$ copies of the nodes in $V=\{v_{1},\dots,v_{n}\}$, labeled as $v_{i}^{j}$, $i\in[n]$, $j\in[2n]$, and two additional nodes denoted by $s$ and $t$. The arcs of $G^{\prime}$ are formed as follows. The vertical arcs connect $v_{i}^{j}$ with $v_{i}^{j+1}$ for $i\in[n]$, $j\in[2n-1]$, node $s$ with $v_{i}^{1}$, $i\in[n]$, and nodes $v_{i}^{2n}$ with $t$, $i\in[n]$. The vertical arcs form $n$ disjoint vertical $s$-$t$ paths $P_{1},\dots,P_{n}$ in $G^{\prime}$, all of which have exactly $12n+$ 1 arcs. The diagonal arcs connect $v_{i}^{2j}$ with $v_{k}^{2j+1}$ for $(v_{i},v_{k})\in A$, $j\in[n-1]$. A sample construction is shown in Figure 1.
Figure 1: A sample digraph $G$ and the corresponding digraph $G^{\prime}$.
Set the budget $\Gamma^{c}=1$, and the recovery parameter $k=2n$ for arc-inclusion (arc-exclusion neighborhood) and $k=4n$ for the arc-symmetric difference neighborhood. All the arcs in $G^{\prime}$ have first-stage costs equal to 0. The second-stage costs of the vertical arcs are in the interval $[0,M]$, and the second-stage costs of the diagonal arcs are in the interval $[M,M]$, where $M$ is a constant such that $M>\Gamma^{c}$. We will show that $G$ has a Hamiltonian path if and only if there is a solution to Rec Rob SP in $G^{\prime}$ with the total cost at most $\frac{\Gamma^{c}}{n}$.
$\Rightarrow$ Let $P$ be a Hamiltonian path that visits the nodes of $G$ in order $v_{i_{1}},\dots,v_{i_{n}}$. Choose a path $X$ in $G^{\prime}$ that contains the arcs
$(s,v_{i_{1}}^{1}),(v_{i_{1}}^{1},v_{i_{1}}^{2}),(v_{i_{1}}^{2},v_{i_{2}}^{3}),(v_{i_{2}}^{3},v_{i_{2}}^{4}),(v_{i_{2}}^{4},v_{i_{3}}^{5}),\dots,(v_{i_{n}}^{2n},t).$
These arcs exist by the construction of $G^{\prime}$ (see Figure 1). The Hamiltonian path in $G$, which visits the nodes of $G$ in order $v_{1},v_{3},v_{2},v_{4}$, corresponds to the $s$-$t$ path composed of the one dashed arc, the dotted arcs, and the one solid arc of the corresponding graph $G^{\prime}$. Observe that $X$ uses at least one arc from each vertical path $P_{1},\dots,P_{n}$. Therefore, each vertical path $P_{i}$ is in $\Phi(X,k)$ and can be used as the second-stage (recovery) path. In the worst second-stage cost scenario for $X$, the budget $\Gamma^{c}$ is shared equally among the arcs of all vertical paths (all vertical paths must have the same second-stage cost). Because the first-stage cost of $X$ is 0 and the second-stage cost of the best second-stage path is $\frac{\Gamma^{c}}{n}$, the total cost of the solution to Rec Rob SP is $\frac{\Gamma^{c}}{n}$.
$\Leftarrow$ Assume that there is a solution to Rec Rob SP with a total cost of at most $\frac{\Gamma^{c}}{n}$. Let $X$ be the first-stage path in this solution. We claim that $X$ must use at least one arc from each vertical path $P_{1},\dots,P_{n}$. Indeed, suppose that one of these paths, say $P_{l}$, does not intersect with $X$. Since $k=2n$, the path $P_{l}\notin\Phi(X,k)$. Since the second-stage costs of the diagonal arcs are equal to $M>\Gamma^{c}$, no diagonal arc can be used in the second-stage path. Therefore, in a worst scenario for $X$, the budget $\Gamma^{c}$ is not allocated to the arcs in path $P_{l}$. Consider scenario $\boldsymbol{c}$ in which $\Gamma^{c}$ is equally shared among all the arcs of the remaining $n-1$ vertical paths. In what follows, every second-stage path in $\Phi(X,k)$ has a cost of at least $\frac{\Gamma^{c}}{n-1}$, which contradicts the assumption that there is a solution to Rec Rob SP with the total cost of at most $\frac{\Gamma^{c}}{n}$. We can now retrieve a Hamiltonian path in $G$ from the diagonal arcs in $X$. Notice first that $X$ has exactly $12n+$ 1 arcs ($G^{\prime}$ is a layered digraph) and contains exactly $n-1$ diagonal arcs. Indeed, the diagonal arcs must connect $n$ disjoint vertical paths, so the number of diagonal arcs is at least $n-1$. On the other hand, the number of diagonal arcs cannot be greater than $n-1$, since the number of arcs in $X$ is equal to $12n+$ 1 (we need $n$ arcs from the vertical paths and two arcs connecting $s$ and $t$ in $X$). Now, it is easy to see that the diagonal arcs in $X$ correspond to the arcs on a Hamiltonian path in $G$. ∎
We now show that the Rec Rob SP problem in general acyclic digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{c})$ is not approximable within $2^{\log^{1-\epsilon}n}$ for any $\epsilon>0$, unless NP $\subseteq$ DTIME$(n^{\mathrm{poly}\log n})$ and it is not approximable within any constant $\alpha>0$ unless P=NP. In our reduction, we will use the following Boolean satisfiability problems with identical inputs, which are known to be strongly NP-hard (see, e.g., [7]):
Input: A 3-CNF formula of $n$ Boolean variables $\{x_{i}\}_{i\in[n]}$ given by $m$ clauses $\{\mathcal{C}_{j}\}_{j\in[m]}$.
3-Sat
Question: Is there a 0-1 assignment to the variables that satisfies all the clauses?
Max 3-Sat
Output: A 0-1 assignment to the variables that maximizes the number of satisfied clauses.
In the following, we will denote by $OPT(\Psi)$ the maximum value of the objective function in a Max 3-Sat instance $\Psi$.
Lemma 1 ([1]).
There is a fixed constant $\gamma>0$ and a polynomial time reduction $\tau$ from 3-Sat to Max 3-Sat such that for every 3-Sat instance $\Psi$:
  1. 1.
    if $\Psi$ is satisfiable then $OPT(\tau(\Psi))=m$,
  2. 2.
    if $\Psi$ is not satisfiable then $OPT(\tau(\Psi))<\frac{m}{1+\gamma}$,
where $m$ is the number of clauses in Max 3-Sat instance $\tau(\Psi)$.
Lemma 1 says equivalently that achieving an approximation ratio of $1+\gamma$ for Max 3-Sat is NP-hard. It has been shown in [9] that such a ratio is of $\frac{8}{7}$ ($\gamma=\frac{1}{7}$).
Lemma 2.
Let $\Psi=(\{x_{i}\}_{i\in[n]},\{\mathcal{C}_{j}\}_{j\in[m]})$ be any instance of Max 3-Sat and $q$ be a positive integer such that $q<2n$ and $\delta$ be any number from $(0,1)$. Then there is a reduction from $\Psi$ to Rec Rob SP instance $\mathcal{I}^{(q)}=(G^{(q)}=(V^{(q)},A^{(q)}),\{C^{(q)}_{e}\}_{e\in A^{(q)}},\mathcal{U}^{(q)}(\Gamma^{c}),\Phi(X,k))$, where $G^{(q)}$ is an acyclic digraph with $|V^{(q)}|\leq n^{O(q)}$, $|A^{(q)}|\leq n^{O(q)}$, $\Phi(X,k)$ is any neighborhood (1)-(3), satisfying the following properties:
  1. 1.
    if $OPT(\Psi)=m$ then $OPT(\mathcal{I}^{(q)})=\frac{1}{m^{q}}$,
  2. 2.
    if $OPT(\Psi)<\delta m$ then $OPT(\mathcal{I}^{(q)})>\frac{1}{\delta^{q}m^{q}}$.
Proof.
Let $\Psi=(\{x_{i}\}_{i\in[n]},\{\mathcal{C}_{j}\}_{j\in[m]})$ be a given an instance of Max 3-Sat. We can assume w.l.o.g. that $x_{i}$ appears in some clause and its negation $\overline{x}_{i}$ appears in some clause. We construct an instance $\mathcal{I}^{(q)}=(G^{(q)}=(V^{(q)},A^{(q)}),\{C^{(q)}_{e}\}_{e\in A^{(q)}},\mathcal{U}^{(q)}(\Gamma^{c}),\Phi(X,k))$ of Rec Rob SP as follows. The underlying acyclic digraph $G^{(q)}$ consists of $m^{q}$ clause gadgets that correspond to the set of all $q$-tuples of the clause set =$\mathfrak{C}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{m}\}$. For every clause tuple $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$, where clause $\mathcal{C}^{j}\in\mathfrak{C}$, $j\in[q]$, has the form $\mathcal{C}_{j}=(l_{1}^{j}\vee l_{2}^{j}\vee l_{3}^{j})$, we build a clause gadget $G^{(q)}_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$, which is a digraph that consists of $2n$ horizontal layers corresponding to variables $x_{i}$ and their negations $\overline{x}_{i}$, $i\in[n]$, and has two distinguished nodes $s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$ and $t_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$. For each $q$-tuple of literals from $\mathcal{L}(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})=\{l_{1}^{1},l_{2}^{1},l_{3}^{1}\}\times\cdots\times\{l_{1}^{q},l_{2}^{q},l_{3}^{q}\}$, which do not contain contradictory literals, we create a literal path from $s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$ to $t_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$. This path contains exactly $q$ literal arcs (solid fat arcs) that correspond to the $q$ components of the literal tuple. The rest of the arcs are dummy arcs (dotted arcs). The path created contains exactly $12n+$ 1 arcs and contains exactly $q$ literal (solid fat) arcs and exactly $2n+1-q$ dummy (dotted arcs). Example are shown in Figure 2a and 3. In Figure 3 the leftmost path from $s_{(\mathcal{C}_{1},\mathcal{C}_{2})}$ to $t_{(\mathcal{C}_{1},\mathcal{C}_{2})}$ corresponds to the tuple of literals $(x_{1},x_{1})$. Observe that two literal arcs are created in the layer corresponding to $x_{1},\overline{x}_{1}$, and a path of length $12n+$ 1 is then created by adding a number of dummy (dotted) arcs. There are at most $3^{q}$ vertical literal paths from $s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$ to $t_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$ in the gadget. They correspond to at most $3^{q}$ $q$-tuples of the literals. from the literal set $\mathcal{L}(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})$. If literal $x_{i}$ (resp. literal $\overline{x}_{i}$), $i\in[n]$, is a component of some literal tuples in $\mathcal{L}(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})$, then we connect the in-literal node $x_{i}$ (resp. the in-literal node $\overline{x}_{i}$), the corresponding literal arcs and the out-literal node $x_{i}$ (resp. the out-literal node $\overline{x}_{i}$) using solid normal arcs to form a path from the in-literal node $x_{i}$ (resp. the in-literal node $\overline{x}_{i}$) to the out-literal node $x_{i}$ (resp. the out-literal node $\overline{x}_{i}$). Otherwise, we add a normal direct arc connecting the in-literal node $x_{i}$ and the out-literal node $x_{i}$ or the in-literal node $\overline{x}_{i}$ and the out-literal node $\overline{x}_{i}$.
Figure 2: (a) A sample of clause gadget $G^{(1)}_{(\mathcal{C}_{1})}$ corresponding to the clause tuple $(\mathcal{C}_{1})$, where $\mathcal{C}_{1}=(x_{1}\vee\overline{x}_{2}\vee x_{4})$. (b) A digraph $G^{(1)}$, that consists of three clause gadgets, corresponding a formula $\Psi=\mathcal{C}_{1}\wedge\mathcal{C}_{2}\wedge\mathcal{C}_{3}$.
Figure 3: A sample clause gadget $G^{(2)}_{(\mathcal{C}_{1},\mathcal{C}_{2})}$, corresponding to the clause tuple $(\mathcal{C}_{1},\mathcal{C}_{2})$, where $\mathcal{C}_{1}=(x_{1}\vee\overline{x}_{2}\vee x_{4})$ and $\mathcal{C}_{2}=(x_{1}\vee x_{2}\vee x_{3})$, includes 8 literal paths. The two leftmost literal paths correspond to the 2-tuples $(x_{1},x_{1})$ and $(\overline{x}_{2},x_{1})$, respectively. The literal path corresponding to $(\overline{x}_{2},x_{2})$ is omitted.
We now construct $G^{(q)}=(V^{(q)},A^{(q)})$ as follows. We add to $G^{(q)}$ all the $m^{q}$ already built clause gadgets $G^{(q)}_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$, where $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$, $n$ variable nodes $u_{1},\dots,u_{n}$ corresponding to variables $x_{1},\dots,x_{n}$, and two distinguished nodes $s$ and $t$. For every $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$ we add two arcs: a dashed arc $(s,s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})})$, a dotted arc $(t_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})},t)$. Next, we add a normal arc $(s,u_{1})$. Then, form two disjoint paths from each variable node $u_{i}$ to the variable node $u_{i+1}$, for $i\in[n-1]$, which traverse through the $m^{q}$ clause gadgets $G^{(q)}_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$, for $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$. This is achieved by adding normal arcs that go to the in-literal nodes $x_{i}$ and $\overline{x}_{i}$ and emanating from the out-literal nodes $x_{i}$ and $\overline{x}_{i}$ of the gadgets. Similarly, for the variable node $u_{n}$, we form two disjoint paths going from $u_{n}$ to the node $t$. These paths are disjoint due to the construction of the clause gadgets. A sample construction of $G^{(1)}=(A^{(1)},V^{(1)})$ is shown in Figure 2b.
The costs (the first and the second stage) of arc $e\in A^{(q)}$ are as follows: $C^{(q)}_{e}=0$ if $e$ is solid (normal or fat), $C^{(q)}_{e}=2$ if $e$ is dotted or dashed; $\hat{c}^{(q)}_{e}=0$ and $\Delta^{(q)}_{e}=0$ if $e$ is fat solid or dotted, $\hat{c}^{(q)}_{e}=2$ and $\Delta^{(q)}_{e}=0$ if $e$ is solid normal, $\hat{c}^{(q)}_{e}=0$ and $\Delta^{(q)}_{e}=1$ if $e$ is dashed. Notice that the uncertainty is only for the costs of the dashed arcs $(s,s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})})$, $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$). Finally, we fix $\Gamma^{c}=1$. We first consider the arc-inclusion neighborhood (see (1)). We set $k=2n+3-q$. Notice that $k>0$, by the assumption that $q<2n$. This completes the reduction. We now show that the reduction produces a gap of $1/\delta^{q}$. The following two claims result from the construction of the instance $\mathcal{I}^{(q)}$.
Claim 1.
There is a one-to-one correspondence between the set of assignments of the formula $\Psi$ and the set of $s$-$t$ paths that use only solid arcs, both normal and fat.
Indeed, consider any 0-1 assignment $\{x_{i}\}_{i\in[n]}$ to the variables. We form a path $X$ as follows: we add arc $(s,u_{1})$ to $X$ and for $i\in[n-1]$ if $x_{i}=1$ then we add the subpath from the variable node $u_{i}$ to the variable node $u_{i+1}$ that traverses the in-literal nodes $x_{i}$ and the out-literal nodes $x_{i}$ of the clause gadgets; otherwise ($x_{i}=0$) we add the subpath from the variable node $u_{i}$ to the variable node $u_{i+1}$ that traverses the in-literal nodes $\overline{x}_{i}$ and the out-literal nodes $\overline{x}_{i}$. Similarly, for the variable node $v_{n}$, we add one of the two disjoint subpaths from the variable node $v_{i}$ to $t$ depending on the value of $x_{i}$. Notice that $X$ is an $s$-$t$ path that uses only solid arcs (both normal and fat). Moreover, for two different 0-1 assignments, the $s$-$t$ paths formed in this way are distinct. It is easily seen that any path $X$ that uses only solid arcs (both normal and fat) determines a 0-1 assignment $\{x_{i}\}_{i\in[n]}$. Namely, if $X$ traverses the in-literal nodes $x_{i}$ and the out-literal nodes $x_{i}$ of the clause gadgets, then we set $x_{i}=1$; otherwise, we set $x_{i}=0$, which proves the claim.
Claim 2.
Let $X$ be any $s$-$t$ path that uses only solid arcs, both normal and fat, and let $\ell$ be the number of clause gadgets that contain at least one path from $s$ to $t$ that intersects exactly $q$ arcs with $X$ ($q$ literal arcs). Then, its total cost (the first and the stage costs) is $\frac{1}{\ell}$, $\frac{1}{m^{q}}\leq\frac{1}{\ell}\leq 1$.
The first-stage cost of $X$ is equal to 0. Any path from $s$ to $t$ that goes through a literal path of a clause gadget intersecting exactly $q$ arcs with $X$ is in the neighborhood $\Phi^{(\mathrm{incl})}(X,2n+3-q)$, so we can replace $X$ with this path in the second stage. Therefore, in a worst second-stage scenario, the budget $\Gamma^{c}$ is shared uniformly among the dashed arcs, arcs $(s,s_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})})$, $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$, that contain these intersecting $s$-$t$ paths. Hence and from the facts that $\Gamma^{c}=1$ and $|\mathfrak{C}^{q}|=m^{q}$, the total cost of $X$ is equal to $\frac{1}{\ell}$, where $\frac{1}{m^{q}}\leq\frac{1}{\ell}\leq 1$.
It remains to prove the implications 1 and 2. Suppose that there exists a 0-1 assignment $\{x_{i}\}_{i\in[n]}$ to the variables that satisfies all the $m$-clauses of the formula $\Psi$, $OPT(\Psi)=m$. By Claim 1, the assignment $\{x_{i}\}_{i\in[n]}$ uniquely determines the first-stage $s$-$t$ path $X$, which uses only solid arcs (both normal and fat) and if $x_{i}=1$, $i\in[n]$, then $X$ uses the arc $(u_{i},x_{i})$; otherwise ($x_{i}=0$) it uses the arc $(u_{i},\overline{x}_{i})$. Since the assignment satisfies $m$ clauses, every clause gadgets $G^{(q)}_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$, $(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})\in\mathfrak{C}^{q}$, contains at least one path from $s$ to $t$, that intersects exactly $q$ literal arcs with $X$, corresponding to a literal tuple in $\mathcal{L}(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})$ satisfying the clauses $\mathcal{C}^{1},\ldots,\mathcal{C}^{q}$. Thus, the total cost of $X$ is $\frac{1}{m^{q}}$, which follows from the fact that $|\mathfrak{C}^{q}|=m^{q}$. Again, by Claim 2, this is a lower bound on the cost of any optimal path. Hence, $X$ is an optimal path and $OPT(\mathcal{I}^{(q)})=\frac{1}{m^{q}}$. This completes the proof of implication 1.
Assume that $OPT(\mathcal{I}^{(q)})\leq\frac{1}{\delta^{q}m^{q}}$. Let $X^{*}$ be an optimal $s$-$t$ path with the cost of $OPT(\mathcal{I}^{(q)})$. Clearly, it uses only solid arcs, both normal and fat, because other arcs have the first-stage cost equal to 2. Let $\ell$ denote the number of clause gadgets that contain at least one path from $s$ to $t$ intersecting exactly $q$ literal arcs from $X^{*}$. These clause gadgets correspond to $\ell$ clause tuples in $\mathfrak{C}^{q}$. By Claim 1, $X^{*}$ uniquely determines the 0-1 assignment $\{x_{i}^{*}\}_{i\in[n]}$ to the variables of $\Psi$. Therefore, in each of the $\ell$ clause gadgets, say, $G^{(q)}_{(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})}$, any $s$-$t$ path intersecting exactly $q$ literal arcs from $X^{*}$ corresponds to a literal tuple in $\mathcal{L}(\mathcal{C}^{1},\ldots,\mathcal{C}^{q})$ that satisfies exactly the $q$ clauses $\mathcal{C}^{1},\ldots,\mathcal{C}^{q}$. From Claim 2 and our assumption, it follows that the total cost of $X^{*}$ is such that $\frac{1}{\ell}\leq\frac{1}{\delta^{q}m^{q}}$ and so $\ell\geq\delta^{q}m^{q}$. Hence, there are at least $\delta^{q}m^{q}$ clause tuples in $\mathfrak{C}^{q}$ such that each clause in the tuple is satisfied under the 0-1 assignment $\{x_{i}^{*}\}_{i\in[n]}$. In consequence, $\{x_{i}^{*}\}_{i\in[n]}$ satisfies at least $\delta m$ clauses of the formula $\Psi$. From the optimality of $X^{*}$ and Claim 1, we conclude that $OPT(\Psi)\geq\delta m$. This proves the implication 2. Accordingly, the optimum value in the instance $\mathcal{I}^{(q)}$ of Rec Rob SP is either $1/m^{q}$ or greater than $1/(\delta m)^{q}$ and so the resulting gap is $1/\delta^{q}$.
To modify the above reduction to obtain the same result for the arc-exclusion and arc-symmetric difference neighborhoods, (2) and(3), we need to ensure that all optimal first-stage paths have the same number of arcs (normal or fat). Choose the arcs of the path from the variable node $u_{i}$ to the variable node $u_{i+1}$ (resp. from $u_{n}$ to $t$) defined for the assignment $x_{i}=1$ or $x_{i}=0$ (resp. $x_{n}=1$ or $x_{n}=0$), see Figure  2b. A raw upper bound on the number of arcs on this path is $r=m^{q}(q3^{q}+1)+1$. We split the first solid arc, $(u_{i},x_{i})$ or $(u_{i},\overline{x}_{i})$, on this path into a number of new solid arcs so that the number of arcs on the path equals $r$. Now, each first-stage path $X$ has exactly $nr+1$ arcs. For the arc-exclusion neighborhood, we set $k=nr+1-q$ and for the arc-symmetric difference, we fix $k=(2n+3)+(nr+1)-2q$. The proofs of implications 1 and 2 for the arc-exclusion and arc-symmetric difference neighborhoods are the same as for the arc-inclusion neighborhood.
It is easily seen that the instance $\mathcal{I}^{(q)}$ has size $N=n^{O(q)}$, where $n$ is the number of variables in the corresponding instance of Max 3-Sat. ∎
Theorem 2.
The Rec Rob SP problem in general acyclic digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{c})$, is not approximable within $2^{\log^{1-\epsilon}N}$ for any $\epsilon>0$, unless NP $\subseteq$ DTIME$(n^{\mathrm{poly}\log n})$, if $k$ is part of the input and $N$ is the input size.
Proof.
The starting point is the gap-preserving reduction from Max 3-Sat to Rec Rob SP (see Lemma 2). By combining this reduction with the reduction from 3-Sat to Max 3-Sat (see Lemma 1) we get a reduction from 3-Sat to Rec Rob SP, which for some $\delta\in(0,1)$ (e.g., $\delta=1/(1+\gamma))$ that ensures:
  1. 1.
    if $\Psi$ is satisfiable then $OPT(\mathcal{I}^{(q)})=\frac{1}{m^{q}}$,
  2. 2.
    if $\Psi$ is not satisfiable then $OPT(\mathcal{I}^{(q)})>\frac{1}{\delta^{q}m^{q}}$.
It has been shown in [9] that $\delta=\frac{7}{8}$.
Let $q=(\log n)^{\beta}$ for some fixed $\beta>0$, where $n$ is the number of variables in an instance of 3-Sat. Now, the size of $\mathcal{I}^{(q)}$ is $N=n^{O(q)}=n^{O(\log^{\beta}n)}=2^{O(\log^{\beta+1}n)}$. Therefore, the running time of the reduction is quasi-polynomial $n^{\text{poly}(\log n)}$. The resulting gap of $\mathcal{I}^{(q)}$ is $\left(\frac{1}{\delta}\right)^{q}=\left(\frac{1}{\delta}\right)^{\log^{\beta}n}=2^{O(\log^{\beta/\beta+1}N)}$.
Assume, on the contrary, that there is a polynomial-time algorithm for Rec Rob SP problem with an approximation ratio of $2^{\log^{1-\epsilon}N}$ for any $\epsilon>1-\frac{\beta}{\beta+1}$. Applying this algorithm to the resulting instance $\mathcal{I}^{(q)}$, one could evaluate whether a formula $\Psi$ in 3-SAT is satisfiable in $O(n^{\mathrm{poly}\log n})$ time. But this would imply NP $\subseteq$ DTIME$(n^{\mathrm{poly}\log n})$, a contradiction. ∎
Corollary 1.
The Rec Rob SP problem in general acyclic digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{c})$, is strongly NP-hard and not approximable within any constant $\alpha>0$, unless P=NP, if $k$ is part of the input.
Proof.
It is sufficient to set $q=\frac{\log\alpha}{\log 1/\delta}$ in the proof of Theorem 2. ∎
We now consider discrete budgeted uncertainty. First, we recall that the Rec Rob SP problem is weakly NP-hard even for a very restricted class of graphs.
Theorem 3 ([8]).
The Rec Rob SP problem is weakly NP-hard in arc series-parallel digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{d})$, if $k$ is part of the input.
In the case of more general graphs, the problem becomes significantly more challenging.
Theorem 4.
The Rec Rob SP problem is strongly NP-hard in layered digraphs with any neighborhood (1)-(3), under $\mathcal{U}(\Gamma^{d})$, and it is also not at all approximable, unless P=NP, if $k$ is part of the input.
Proof.
Given an instance $G$ of Hamiltonian Path, the construction of $G^{\prime}$ is the same as in the proof of Theorem 1. We again set the recovery parameter $k=2n$ for the arc-inclusion (arc-exclusion) neighborhood and $k=4n$ for the arc-symmetric difference neighborhood. The first-stage costs of all arcs in $G^{\prime}$ are 0. The second-stage costs of all vertical arcs in $G^{\prime}$ are in the interval $[0,0]$, except for the dashed arcs $(s,v_{i}^{1})$, $i\in[n]$, whose second-stage costs are in the interval $[0,1]$ (see the dashed arcs in Figure 1). The second-stage costs of all diagonal arcs are in the interval $[1,1]$. Let $\Gamma^{d}=n-1$. We claim that there is a Hamiltonian path in $G$ if and only if there is a solution to Rec Rob SP in $G^{\prime}$ with the cost equal to 0. The proof of this claim is based on the same reasoning as in the proof of Theorem 1. Graph $G$ has a Hamiltonian path if and only if there is a first-stage path $X$ in $G^{\prime}$ that uses at least one arc from each vertical path $P_{1},\dots,P_{n}$. Because, in the worst second-stage cost scenario, we can set the cost of at most $n-1$ dashed arcs to 1, $G$ has a Hamiltonian path if and only if there is a second-stage path in $\Phi(X,k)$ with the cost equal to 0. Using the fact that the first-stage cost of each solution equals 0, we are done. ∎
Acknowledgements
The authors were supported by the National Science Centre, Poland, grant 2022/45/B/HS4/00355.